OpenStudy (anonymous):

Find the two linearly independent solution of d^2y/dx^2 + x dy/dx + y =0 when the series expansion is about the origin ..... Please I will give a medal

OpenStudy (perl):

$$ \Large y(x) =a_0e^{\frac{-x^2}{2}} +a_1\sum_{n=0}^{\infty}\frac{(-1)^nx^{2n+1}(2n+1)!}{2^n n!} $$

OpenStudy (perl):

ill get it started, some of the steps

OpenStudy (perl):

$$ \text{we want to solve } \\ \Large y ' ' + x y' + y = 0 \\ \text {Assume your solution is } \\ \Large y(x)=\sum_{n=0}^{\infty}a_n x^n\\ \text {then it follows that} \\ \Large y'(x) = \sum_{n=1}^{\infty}n \cdot a_n x^{n-1} \\ \Large y''(x) = \sum_{n=2}^{\infty}n(n-1) \cdot a_n x^{n-2}\\ \text {Substitute: }\Large y ' ' + x y' + y = 0 \\ \large \sum_{n=2}^{\infty}n(n-1) \cdot a_n x^{n-2} + x\sum_{n=1}^{\infty}n \cdot a_n x^{n-1} + \sum_{n=0}^{\infty}a_n x^n=0\\ \large \iff \\ \large \sum_{n=2}^{\infty}n(n-1) \cdot a_n x^{n-2} + \sum_{n=1}^{\infty}n \cdot a_n x^{n} + \sum_{n=0}^{\infty}a_n x^n=0\\ \large \iff \\ \large \sum_{n=0}^{\infty}(n+2)(n+1) \cdot a_{n+2} x^{n} + \sum_{n=1}^{\infty}n \cdot a_n x^{n} + \sum_{n=0}^{\infty}a_{n+2} x^n=0\\ \large \iff\\ \large \sum_{n=0}^{\infty}(n+2)(n+1) \cdot a_{n+2} x^{n} + \sum_{n=1}^{\infty}n \cdot a_n x^{n} + \sum_{n=0}^{\infty}a_n x^n=0\\ $$

OpenStudy (anonymous):

thanks so much i understood perfectly. but please can you help me find the initial value problem of that equation when y(0)=1 and dy/dx(0)=0 god bless